# A YTD correspondence for constant scalar curvature metrics

### v1

@@ 2 Preliminaries
Based on this simple fact, a remarkable result was obtained by Li [61]: if $\{u_{t}\}\in\mathcal{R}^{1}$ is a geodesic ray with $Ent\{u_{t}\}<\infty$, then it is maximal in the sense of [10, Definition 6.5]. More specifically, we have that (cf. [10, Lemma 5.7], [75, Theorem 4.4.1])
$\{u_{t}^{\mathcal{T}_{m}}\}\xrightarrow{d_{1}^{c}}\{u_{t}\}\text{ as }m\to\infty.$
@@ 3 Transcendental quantization
Let $(X,\omega)$ be a compact Kähler manifold of dimension $n$. Recall that for any $\beta>0$ and $u\in\mathcal{E}^{1}$, $e^{-\beta u}\in L^{p}$ for any $p>1$ [94]. Thus, by the theory of Monge-Ampère equations [9, Theorem C], there exists a unique continuous $\omega$-psh function $u^{\beta}$ satisfying
@@
Based on the speculations of [6, Section 3.1], we call $u^{\beta}$ the transcendental quantization of $u$ at level $\beta$ (see [34, Appendix A.3] for a short survey on this topic).
@@
The first point is trivial and the second follows from the comparison principle [6, Lemma 2.5] (whose proof applies to potentials of $\mathcal{E}^{1}$). Finally, we argue the last point:
@@
We first show that $u^{\beta}\xrightarrow{a.e.}u$ as $\beta\to\infty$. Let $u\in\mathcal{E}^{1}$ and $\delta>1$. Consider $u_{\delta}:=\delta^{-1}P(\delta u)\in\mathcal{E}^{1}$ as in [38, Proposition 2.15] (Proposition 2.18 in the published version). Then $u_{\delta}\leq u\text{ and }\omega+\mathrm{dd^{c}}u_{\delta}\geq(\frac{\delta-1}{\delta})\omega.$ Thus:
@@
Applying [6, Lemma 2.5] again, we obtain that
@@
In particular, $\liminf_{\beta\to\infty}u^{\beta}(x)\geq u_{\delta}(x),\ \text{for all }\delta>1,x\in X.$ Since $u_{\delta}\xrightarrow{a.e.}u$ as $\delta\searrow 1$ ([38, Proposition 2.15]), we obtain that
@@
We now bound $u^{\beta}$ from above. Pick a smooth sequence $u_{k}\searrow u$ pointwise. Using the fact that $u_{k}^{\beta}\geq u^{\beta}$ and that $u_{k}^{\beta}\xrightarrow{C^{0}}u_{k}$ as $\beta\to\infty$ ([31, Theorem 1.1], [34, Theorem A.10]),
@@
Sending $\beta\to\infty$, using dominated convergence and [32, Theorem 3],
@@
Finally, sending $\delta\searrow 1$ and using [38, Proposition 2.15], we conclude. ∎
@@
Thus $Ent(u)\geq Ent(u^{\beta})$ for any $u\in\mathcal{E}^{1}$. So we conclude since $Ent(\cdot)$ is $d_{1}$-lsc ([8], cf. [34, Corollary 4.40]). ∎
@@
For any $u\in\mathcal{E}^{1}$,
$Ent(u)\geq\beta(I(u^{\beta})-I(u))\geq Ent(u^{\beta}).$
The next result says that our quantization procedure improves convergence:
@@
Assume that $u_{i}\xrightarrow{L^{1}}u$ in $\mathcal{E}^{1}$, then $u^{\beta}_{i}-u^{\beta}||_{C^{0}}\to 0$ for any $\beta>0$. In particular, $u_{i}^{\beta}\xrightarrow{d_{1}}u^{\beta}$.
@@
Since for any $p>0$, $e^{-pu_{i}}$ and $e^{-pu}$ are integrable by [94], [45, Theorem 0.2] implies that $\int_{X}|e^{-pu_{i}}-e^{-pu}|\omega^{n}\to 0.$ In particular, there exists $A_{p}>0$ such that
@@
On the other hand, by the stability estimate [55, Remark 3.2], for any $\beta>0$,
@@
For any $u\in\mathcal{E}^{1}$ and $\beta>0$,
$Ent^{\beta}(u)=\beta(I(u^{\beta})-I(u)).$
@@
That $Ent^{\beta}(u)\geq\beta(I(u^{\beta})-I(u))$ follows from (18) and (16). Conversely, for $v\in\mathcal{E}^{1}_{\omega}$ we have
@@
The first point follow from Lemma 3.5 and that $I(\cdot)$ is $L^{1}$-usc [8] (see also [34, Corollary 4.14]). Then second follows from Lemma 3.5 as well. ∎
@@
Let $0<\beta_{1}\leq\beta_{1}$. Then for any $u\in\mathcal{E}^{1}$, $Ent^{\beta_{1}}(u)\leq Ent^{\beta_{2}}(u).$
@@
To this end, we differentiate (16) with respect to $\beta$:
$\Delta_{\omega_{u^{\beta}}}\frac{du^{\beta}}{d\beta}=(u^{\beta}-u)+\beta\frac{du^{\beta}}{d\beta}.$
Integrating against $\omega^{n}_{u^{\beta}}$, we derive that $\frac{\beta}{V}\int_{X}\frac{du^{\beta}}{d\beta}\omega^{n}_{u^{\beta}}=-\frac{1}{V}\int_{X}(u^{\beta}-u)\omega^{n}_{u^{\beta}}.$ Thus,
@@
For $u\in\mathcal{E}^{1}$ we have $Ent(u^{\beta})\leq Ent^{\beta}(u)\leq Ent(u)$ and $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
@@
In the classical setting of algebraic quantization, a useful positivity result due to Berndtsson roughly says that the quantization of a subgeodesic is still a subgeodesic [13, Proposition 3.1], a crucial point in the second author’s alternative proof of the YTD conjecture for Kähler-Einstein type metrics [96]. Later in §7, a transcendental replacement for Berndtsson’s result will be needed. It is given by the following positivity result that can be extracted from Păun’s paper [74, §3] (cf. also [29, Theorem 4.1]).
@@
we obtain a smooth semipositive $(1,1)$-form on $X\times\Sigma$ whose restriction on each fiber $X\times\{z\}$ is the Kähler form $\beta\omega_{u_{t}}+\operatorname{Ric}(\omega)$. Let
$\Phi_{z}:=\beta u_{-\log|z|^{2}}^{\beta}-\beta u_{-\log|z|^{2}}.$
Then the Monge–Ampère equation of $u_{t}^{\beta}$ (16) translates to
@@
To this end, we need to further regularize $u_{t,\varepsilon}$. For any $0<\delta<\varepsilon$, by [15, Theorem 2] we can find smooth approximating functions $U_{\varepsilon,\delta}$ on $X\times\{e^{-1}+\delta<|z|^{2}<1-\delta\}$ such that
@@
Therefore, it suffices that argue that $\{u^{\beta}_{t,\varepsilon,\delta}\}$ is subgeodesic for all $\beta>2A$, which now follows Theorem 3.13, and the discussion following it. ∎
@@
If $\operatorname{Ric}(\omega)\geq 0$, then we can take $\eta=0$ (using a limit process $\eta\searrow 0$) in the above result. So in this case $t\mapsto u_{t}$ is subgeodesic for any $\beta>0$ whenever $t\mapsto u_{t}$ is.
@@ 4 Properness of the $K^{\beta}$ energy
In our first technical result we show that the distance $d_{1}(u,u^{\beta})$ is comparable to the growth of $\frac{1}{\beta}Ent^{\beta}(u)$. The role of this result in our work is analogous with the role of the partial $C^{0}$ estimate from the Fano case, first proposed by Tian (cf. [84, (0.2)]).
@@
We control $d_{1}(P(\delta u),u)$ as follows (using [32, Theorem 3] and [35, Lemma 4.4]):
@@
We bound the terms on the last line one by one. By [10, Lemma A.2], we have
@@
where $M:=\max\{\mathcal{I}(u),\mathcal{I}(u^{\beta})\},$ and $\mathcal{I}(u):=\frac{1}{V}\int_{X}u(\omega^{n}-\omega^{n}_{u})$ denotes Aubin’s $\mathcal{I}$-functional. We assumed that $u\leq 0$, so $u^{\beta}\leq 0$ as well. Due to [32, Theorem 3],
@@
First, it is well known that (see e.g. [34, Lemma 3.45]) $|\sup_{X}u-\frac{1}{V}\int_{X}u\omega^{n}|\leq C.$ Therefore, it suffices to argue that
@@
Using [10, Lemma A.2] exactly as in the proof of the previous argument we obtain that
$\int_{X}(u^{\beta}-u)\omega^{n}\leq Cd_{1}(u^{\beta},u)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}.$
By (4) we conclude (23). ∎
@@
because $d_{1}(0,u)=-I(u)=J(u)-\int_{X}u\omega^{n},$ and it is well known that $\sup_{X}u-\int_{X}u\omega^{n}$ is always uniformly bounded (cf. [39, Proposition 5.5]).
@@
Finally, using that $|\mathcal{J}_{\operatorname{Ric}(\omega)}(u)|\leq Cd_{1}(0,u)$ ([36, Proposition 2.5]) and choosing $\beta$ to be sufficiently large (depending only on $X,\omega,n,\delta$), we obtain $\delta^{\prime}>0$ and $C_{0}^{\prime}>0$ such that $K^{\beta}(u)\geq\delta^{\prime}d_{1}(0,u)-C_{0}^{\prime}$ for any $u\in\mathcal{E}^{1}$ with $\sup_{X}u=0$. This completes the proof. ∎

### v2

@@ 2 Preliminaries
Using this simple fact, a remarkable result was obtained by Li [70]: if $\{u_{t}\}\in\mathcal{R}^{1}$ is a geodesic ray with $Ent\{u_{t}\}<\infty$, then it is maximal in the sense of [9, Definition 6.5]. More specifically, we have $d_{1}(\{u_{t}^{\mathcal{T}_{m}}\},\{u_{t}\})\to 0\text{ as }m\to\infty$ (cf. [9, Lemma 5.7], [85, Theorem 4.4.1]).
Models and non-Archimedean geometry.
In this work we also need to work with models, introduced in [70, Definition 2.7]. These are generalizations of test configurations $(\mathcal{X},\mathcal{L})$, where $\mathcal{L}$ is only relatively big on $\mathcal{X}$ (namely, $\mathcal{L}+c\mathcal{X}_{0}$ is big for some $c>0$). One can also naturally attach a geodesic ray to a model, since a model determines a filtration (as shown in [70, Definition 2.7]), and a filtration in turn also determines a geodesic ray [86, §9], which is constructed as an increasing limit of a sequence of geodesic rays of test configurations. As a consequence, the singularity type of such a geodesic ray is typically no longer of algebraic type (c.f (21)), but instead shall be seen as certain envelope of algebraic singularities. A convenient way of describing it is to use some terms from non-Archimedean geometry, that we now turn to explain.
Let $X$ be a projective manifold. Following [24], denote by $X^{an}$ the Berkovich analytification of $X$ with respect to the trivial absolute value on the ground field $\mathbb{C}$}. We will mostly work with the subspace $X_{\mathbb{Q}}^{div}$ consisting of $\mathbb{Q}$-valued divisorial valuations on $X$. Elements in $X^{div}_{\mathbb{Q}}$ can be explicitly described as follows. Any $w\in X^{div}_{\mathbb{Q}}$ is of the form $\lambda\mathrm{ord}_{F}$, with $\lambda\in\mathbb{Q}_{\geq 0}$ and $F\subset Y\xrightarrow{\tau}X$ being some prime divisor over $X$. When $\lambda=0$, we get the trivial valuation $v_{triv}.$ Moreover, we define the log discrepancy of such $w$ to be
$A_{X}(w)=A_{X}(\lambda\mathrm{ord}_{F}):=\lambda(1+\mathrm{ord}_{F}(K_{Y}-\tau^{*}K_{X})).$
The notion of log discrepancy further extends in a linear way to the more general quasi-monomial valuations in $X^{an}$ (see e.g. [101, Definition 1.34]).
Let $(X,L)$ be a polarized Kähler manifold. Let $\mathcal{H}^{NA}$ be the space of non-Archimedean potentials coming from test configurations of $(X,L)$ (see also [9, §3.2]). Any element of $\mathcal{H}^{NA}$ is of the form $u^{\mathcal{T}}$, where $\mathcal{T}$ is some test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$ of $(X,L)$. Here $u^{\mathcal{T}}$ is a continuous function on $X^{an}$, whose value on any $w\in X^{an}$ is explicitly determined by
$u^{\mathcal{T}}(w):=-\sigma(w)(D),$ (23)
where $\sigma(w)$ denotes the Gauss extension of $w$ (see [21, Definition 4.3]) and $D$ is given by (20). When $w\in X^{div}_{\mathbb{Q}}$, it follows from [21, Theorem 4.6] that, up to passing to a higher model, $\sigma(w)$ is of the form $\frac{1}{b_{E}}\mathrm{ord}_{E}$ for some irreducible component $E$ in $\mathcal{X}_{0}$, with $b_{E}:=\mathrm{ord}_{E}(\mathcal{X}_{0})$. Moreover, by [25, Theorem A.10] we have the relation $A_{X}(w)=\frac{1}{b_{E}}A_{X\times{\mathbb{P}}^{1}}(E)-1.$ If we assume in addition that the geodesic ray $\{u_{t}^{\mathcal{T}}\}$ attached to $\mathcal{T}$ satisfies $u_{t}^{\mathcal{T}}\leq 0$, then $u^{\mathcal{T}}$ can be alternatively defined using Lelong numbers (see [9, Definition 4.2]):
$u^{\mathcal{T}}(w)=-\frac{1}{b_{E}}\nu(u^{\mathcal{T}}_{-\log|z|^{2}},E).$ (24)
Here recall that $\nu(\phi,E)$ denotes the Lelong number of a qpsh function $\phi$ along $E$, which is defined as
Denote by $PSH^{NA}$ the space of $L$-psh functions on $X^{an}$. Recall that $u\in PSH^{NA}$ if and only if $u\not\equiv$-$\infty$ and $u$ is equal to the decreasing limit of a sequence of functions in $\mathcal{H}^{NA}$.
Now let $\mathcal{M}:=(\mathcal{X},\mathcal{L})$ be a model of $(X,L)$, so $\mathcal{L}$ is not assumed to be relatively semi-ample. In this case, we can still define $D$ as in (20). Put
$f_{D}(w):=-\sigma(w)(D),\ w\in X^{an}.$
Then $f_{D}$ is a continuous function on $X^{an}$, but in general $f_{D}(w)\notin\mathcal{H}^{NA}$. By [20, Theorem 8.3] the non-Archimedean potential $u^{\mathcal{M}}$ associated to the model is then given by the following envelope:
$u^{\mathcal{M}}:=P(f_{D}):=\sup\{v\in PSH^{NA}:v\leq f_{D}\}.$ (25)
It follows from [20, Theorem 8.5] (see also [23, Proposition 4.5]) that $u^{\mathcal{M}}$ is continuous.
We can also describe $u^{\mathcal{M}}$ using Lelong numbers as before. Let $\{u_{t}^{\mathcal{M}}\}$ be the geodesic ray attached to $\mathcal{M}$. After subtracting some multiple of $\mathcal{X}_{0}$ from $\mathcal{L}$ we can always assume that $u_{t}^{\mathcal{M}}\leq 0$. Then for any $w\in X^{div}_{\mathbb{Q}}$ with $\sigma(w)=\frac{1}{b_{E}}\mathrm{ord}_{E}$, we have
$u^{\mathcal{M}}(w)=-\frac{1}{b_{E}}\nu(u^{\mathcal{M}}_{-\log|z|^{2}},E).$ (26)
This characterization of the singularity types of model rays will be used in §9.
Complex reductive Lie groups.
In the proof of Theorem 1.6, some classical Lie group theory will be needed, which we now briefly recall. Let $\mathbb{G}$} be a complex reductive Lie group with maximal compact subgroup $\mathbb{K}$}. Due to reductivity of $\mathbb{G}$} there exists a compact semi-simple group $\mathbb{L}\subset{\mathbb{K}}$ such that $\mathbb{L}_{\mathbb{C}}\leq{\mathbb{G}}$ and
${\mathbb{G}}\simeq\big(\mathbb{L}_{\mathbb{C}}\times Z({\mathbb{G}})\big)/\Gamma_{1},$ (27)
where $\Gamma_{1}$ is a discrete subgroup of $\mathbb{L}_{\mathbb{C}}\times Z(G)$, and $Z({\mathbb{G}})$ is the center of $\mathbb{G}$} that can be further written as
$Z({\mathbb{G}})\simeq({\mathbb{C}}^{*})^{m}\times({\mathbb{C}}^{k}/\Lambda),$ (28)
where $\Lambda\leq{\mathbb{C}}^{k}$ is a sublattice making ${\mathbb{C}}^{k}/\Lambda\leq{\mathbb{K}}$ compact [16, Proposition, p. 181]. Using our notation, $Z({\mathbb{G}})\cap{\mathbb{K}}=(S^{1})^{m}\times({\mathbb{C}}^{k}/\Lambda)$ is the maximal compact subgroup of $Z({\mathbb{G}})$ and we have the following decomposition of $\mathbb{K}$}:
${\mathbb{K}}\simeq\big(\mathbb{L}\times(S^{1})^{m}\times{\mathbb{C}}^{k}/\Lambda\big)/\Gamma_{2},$ (29)
where $\Gamma_{2}$ is a discrete subgroup of $\mathbb{L}\times(S^{1})^{m}\times{\mathbb{C}}^{k}/\Lambda$.
Finally, we denote by ${\mathbb{T}}=({\mathbb{C}}^{*})^{m}\leq Z({\mathbb{G}})$ the maximal torus of $Z({\mathbb{G}})$, satisfying ${\mathbb{K}}\cap{\mathbb{T}}=(S^{1})^{m}$. With the above decompositions in hand, let us recall the following well known result relating the normalizer $N_{\mathbb{G}}({\mathbb{K}})$ of the maximal compact group $\mathbb{K}$} to the center of $\mathbb{G}$} and $\mathbb{T}$} [69, Proposition B.1]:
Lemma 2.1.
Let $\mathbb{G}$} be a complex reductive Lie group with maximal compact subgroup $\mathbb{K}$}. With the above notation, $N_{\mathbb{G}}({\mathbb{K}})=Z({\mathbb{G}}){\mathbb{K}}={\mathbb{T}}{\mathbb{K}}$.
@@ 3 Transcendental quantization
Let $(X,\omega)$ be a compact Kähler manifold of dimension $n$. Recall that for any $\beta>0$ and $u\in\mathcal{E}^{1}$, $e^{-\beta u}\in L^{p}$ for any $p>1$ [105]. Thus, by the theory of Monge–Ampère equations [8, Theorem C], there exists a unique continuous $\omega$-psh function $u^{\beta}$ satisfying
@@
Based on the speculations of [5, Section 3.1], we call $u^{\beta}$ the transcendental quantization of $u$ at level $\beta$ (see [38, Appendix A.3] for a short survey on this topic).
@@
The first point is trivial and the second follows from the comparison principle [5, Lemma 2.5] (whose proof applies to potentials of $\mathcal{E}^{1}$). Finally, we argue the last point:
@@
We first show that $u^{\beta}\xrightarrow{a.e.}u$ as $\beta\to\infty$. Let $u\in\mathcal{E}^{1}$ and $\delta>1$. Consider $u_{\delta}:=\delta^{-1}P(\delta u)\in\mathcal{E}^{1}$ as in [42, Proposition 2.15] (Proposition 2.18 in the published version). Then $u_{\delta}\leq u\text{ and }\omega+\mathrm{dd^{c}}u_{\delta}\geq(\frac{\delta-1}{\delta})\omega.$ Thus:
@@
Applying [5, Lemma 2.5] again, we obtain that
@@
In particular, $\liminf_{\beta\to\infty}u^{\beta}(x)\geq u_{\delta}(x),\ \text{for all }\delta>1,x\in X.$ Since $u_{\delta}\xrightarrow{a.e.}u$ as $\delta\searrow 1$ ([42, Proposition 2.15]), we obtain that
@@
We now bound $u^{\beta}$ from above. Pick a smooth sequence $u_{k}\searrow u$ pointwise. Using the fact that $u_{k}^{\beta}\geq u^{\beta}$ and that $u_{k}^{\beta}\xrightarrow{C^{0}}u_{k}$ as $\beta\to\infty$ ([35, Theorem 1.1], [38, Theorem A.10]),
@@
Sending $\beta\to\infty$, using dominated convergence and [36, Theorem 3],
@@
Finally, sending $\delta\searrow 1$ and using [42, Proposition 2.15], we conclude. ∎
@@
Thus $Ent(u)\geq Ent(u^{\beta})$ for any $u\in\mathcal{E}^{1}$. So we conclude since $Ent(\cdot)$ is $d_{1}$-lsc ([7], cf. [38, Corollary 4.40]). ∎
@@
For any $u\in\mathcal{E}^{1}$, $Ent(u)\geq\beta(I(u^{\beta})-I(u))\geq Ent(u^{\beta}).$
The next result shows that our quantization procedure improves convergence:
@@
If $u_{i}\xrightarrow{L^{1}}u$ in $\mathcal{E}^{1}$, then $u^{\beta}_{i}-u^{\beta}||_{C^{0}}\to 0$ for any $\beta>0$. In particular, $u_{i}^{\beta}\xrightarrow{d_{1}}u^{\beta}$.
@@
Since for any $p>0$, $e^{-pu_{i}}$ and $e^{-pu}$ are integrable by [105], [49, Theorem 0.2] implies that $\int_{X}|e^{-pu_{i}}-e^{-pu}|\omega^{n}\to 0.$ In particular, there exists $A_{p}>0$ such that
@@
On the other hand, by the stability estimate [62, Remark 3.2], for any $\beta>0$,
@@
For any $u\in\mathcal{E}^{1}$ and $\beta>0$, $Ent^{\beta}(u)=\beta(I(u^{\beta})-I(u)).$
@@
That $Ent^{\beta}(u)\geq\beta(I(u^{\beta})-I(u))$ follows from (32) and (30). Conversely, for $v\in\mathcal{E}^{1}_{\omega}$ Jensen’s inequality implies that
@@
The first point follow from Lemma 3.5 and that $I(\cdot)$ is $L^{1}$-usc [7] (see also [38, Corollary 4.14]). Then second follows from Lemma 3.5 as well. ∎
@@
Let $0<\beta_{1}\leq\beta_{2}$. Then for any $u\in\mathcal{E}^{1}$, $Ent^{\beta_{1}}(u)\leq Ent^{\beta_{2}}(u).$
@@
To this end, we differentiate (30) with respect to $\beta$ to get $\Delta_{\omega_{u^{\beta}}}\frac{du^{\beta}}{d\beta}=(u^{\beta}-u)+\beta\frac{du^{\beta}}{d\beta}.$ Integrating against $\omega^{n}_{u^{\beta}}$, we derive that $\frac{\beta}{V}\int_{X}\frac{du^{\beta}}{d\beta}\omega^{n}_{u^{\beta}}=-\frac{1}{V}\int_{X}(u^{\beta}-u)\omega^{n}_{u^{\beta}}.$ Thus,
@@
For $u\in\mathcal{E}^{1}$ we have $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
@@
In the classical setting of algebraic quantization, a useful positivity result due to Berndtsson roughly says that the quantization of a subgeodesic is still a subgeodesic [12, Proposition 3.1], a crucial point in the second author’s alternative proof of the YTD conjecture for Kähler-Einstein type metrics [107]. Later in §6, a transcendental replacement for Berndtsson’s result will be needed. It is given by the following positivity result that can be extracted from Păun’s paper [84, §3] (cf. also [33, Theorem 4.1]).
@@
we obtain a smooth semipositive $(1,1)$-form on $X\times\Sigma$ whose restriction on each fiber $X\times\{z\}$ is the Kähler form $\beta\omega_{u_{t}}+\operatorname{Ric}(\omega)$. Let $\Phi_{z}:=\beta u_{-\log|z|^{2}}^{\beta}-\beta u_{-\log|z|^{2}}.$ Then the Monge–Ampère equation of $u_{t}^{\beta}$ (30) translates to
@@
To this end, we need to further regularize $u_{t,\varepsilon}$. For any $0<\delta<\varepsilon$, by [14, Theorem 2] we can find smooth approximating functions $U_{\varepsilon,\delta}$ on $X\times\{e^{-1}+\delta<|z|^{2}<1-\delta\}$ such that
@@
Therefore, it suffices to argue that $\{u^{\beta}_{t,\varepsilon,\delta}\}$ is subgeodesic for all $\beta>2A$, which now follows Theorem 3.13, and the discussion following it. ∎
@@
If $\operatorname{Ric}(\omega)\geq 0$, then we can take $\eta=0$ (using a limit process $\eta\searrow 0$) in the above result. So in this case $t\mapsto u^{\beta}_{t}$ is subgeodesic for any $\beta>0$ whenever $t\mapsto u_{t}$ is.
@@ 4 Properness of the $K^{\beta}$ energy
In our first technical result we show that the distance $d_{1}(u,u^{\beta})$ is comparable to the growth of $\frac{1}{\beta}Ent^{\beta}(u)$. The role of this result in our work is analogous with the role of the partial $C^{0}$ estimate from the Fano case, first proposed by Tian (cf. [94, (0.2)]).
@@
We control $d_{1}(P(\delta u),u)$ as follows (using [36, Theorem 3] and [39, Lemma 4.4]):
@@
We bound the terms on the last line one by one. By [9, Lemma A.2], we have
@@
where $M:=\max\{\mathcal{I}(u),\mathcal{I}(u^{\beta})\},$ and $\mathcal{I}(u):=\frac{1}{V}\int_{X}u(\omega^{n}-\omega^{n}_{u})$ denotes Aubin’s $\mathcal{I}$-functional. We assumed that $u\leq 0$, so $u^{\beta}\leq 0$ as well. Due to [36, Theorem 3],
@@
First, it is well known that (see e.g. [38, Lemma 3.45]) $|\sup_{X}u-\frac{1}{V}\int_{X}u\omega^{n}|\leq C.$ Therefore, it suffices to argue that
@@
Using [9, Lemma A.2] exactly as in the proof of the previous argument we obtain that $\Big|\int_{X}(u^{\beta}-u)\omega^{n}\Big|\leq Cd_{1}(u^{\beta},u)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}.$ By (4) we conclude (37). ∎
@@
because $d_{1}(0,u)=-I(u)=J(u)-\int_{X}u\omega^{n},$ and it is well known that $\sup_{X}u-\int_{X}u\omega^{n}$ is always uniformly bounded (cf. [43, Proposition 5.5]).
@@
Finally, using that $|\mathcal{J}_{\operatorname{Ric}(\omega)}(u)|\leq Cd_{1}(0,u)$ ([40, Proposition 2.5]) and choosing $\beta$ to be sufficiently large (depending only on $X,\omega,n,\delta$), we obtain $\delta^{\prime}>0$ and $C_{0}^{\prime}>0$ such that $K^{\beta}(u)\geq\delta^{\prime}d_{1}(0,u)-C_{0}^{\prime}$ for any $u\in\mathcal{E}^{1}$ with $\sup_{X}u=0$. This completes the proof. ∎